OpenStudy (anonymous):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)? A) x2 + y2 − 4x + 2y + 1 = 0 B) x2 + y2 + 4x − 2y + 1 = 0 C) x2 + y2 + 4x − 2y + 9 = 0 D) x2 − y2 + 2x + y + 1 = 0 which answer choice ?

OpenStudy (compassionate):

For a circle who's center is (h, k), the general form is \[(x - h)^2 + (y - k)^2 = r^2\]

OpenStudy (campbell_st):

well the distance from (-2, 1) to (-4, 1) is the radius the general form is \[(x- h)^2 + (y - k)^2 = r^2\] where (h, k) is the centre and r is the radius... so find the radius and substitute your centre and radius values

OpenStudy (anonymous):

thank you but I don't know how to substitute my centre and radius values

OpenStudy (campbell_st):

so what is your radius value..?

OpenStudy (anonymous):

I'm not sure

OpenStudy (campbell_st):

ok... so this may help 

OpenStudy (campbell_st):